40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5207     Entry time: Fri Aug 12 15:16:56 2011
Author: jamie 
Type: Update 
Category: SUS 
Subject: today's SUS overview 

Here's an update of the suspensions, after yesterdays in-vacuum work and OSEM tweaking:

  • PRM and ETMY are completely messed up.  The spectra are so bad that I'm not going to bother posting anything.   ETMY has extreme sensor voltages that indicate that it's maybe stuck to one of the OSEMS.  PRM voltages look nominal, so I have no idea what's going on there.
  • ITMY is much improved, but it could still use some work
  • SRM is a little worse than what it was yesterday, but we've done a lot of work on the ITMY table, such as moving ITMY suspension and rebalancing the table.
  • BS looks for some reason slightly better than it did yesterday
TM   M cond(B)
SRM SRM.png

      pit     yaw     pos     side    butt
UL    0.828   1.041   1.142  -0.135   1.057 
UR    1.061  -0.959   1.081  -0.063  -1.058 
LR   -0.939  -0.956   0.858  -0.036   0.849 
LL   -1.172   1.044   0.919  -0.108  -1.035 
SD    0.196  -0.024   1.861   1.000   0.043

4.20951
ITMY  ITMY.png       pit     yaw     pos     side    butt
UL    1.141   0.177   1.193  -0.058   0.922 
UR    0.052  -1.823   0.766  -0.031  -0.974 
LR   -1.948  -0.082   0.807  -0.013   1.147 
LL   -0.859   1.918   1.234  -0.040  -0.957 
SD   -1.916   2.178   3.558   1.000   0.635 
7.70705
BS  BS.png       pit     yaw     pos     side    butt
UL    1.589   0.694   0.182   0.302   1.042 
UR    0.157  -1.306   1.842  -0.176  -0.963 
LR   -1.843  -0.322   1.818  -0.213   0.957 
LL   -0.411   1.678   0.158   0.265  -1.038 
SD    0.754   0.298  -3.142   1.000   0.053
6.12779

 

ELOG V3.1.3-